OpenStudy (anonymous):

I have a question: Does square root(o(square)+3(square) = 0+3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Faris, in the equation editor, type sqrt{} to do a square root, and ^{} to do powers.

OpenStudy (asnaseer):

no:\[\sqrt{a^2+b^2}\ne a+b\]\[\sqrt{0^2+3^2}=\sqrt{0+9}=\sqrt{9}=\pm3\]

OpenStudy (asnaseer):

e.g.:\[\sqrt{4^2+3^2}=\sqrt{16+9}=\sqrt{25}=\pm5\]\[\sqrt{4^2+3^2}\ne4+3\]

OpenStudy (anonymous):

Good job @asnaseer!! Sorry @Faris_Waleed , I was wrong.

OpenStudy (asnaseer):

thx Josh :)

OpenStudy (asnaseer):

and remember - we all make mistakes - that's what makes us human :)